a) The total population of a region in a certain country is 15,500. A village with a population of 620 is situated in the region. Express the population of the village as a percentage of the total population of the region.
b) Another village in the region has a population of 992. If a grant of $ 65,000 is to be shared between the two villages in proportion to their population, calculate how much each village receives.

Answer by rwm(914)

a)
620/15500=.04 or 4 %
This was asked many years ago question 155011
http://www.algebra.com/algebra/homework/word/evaluation/Evaluation_Word_Problems.faq.question.155011.html
b)
620+992=population of the two villages=1612
620/1612 *65000=25000 portion to village one
992/1612*65000=40000 portion to village two